OpenStudy (anonymous):

If the discriminant is 0, how many real solutions does the quadratic equation have? 2 1 3 0

OpenStudy (anonymous):

I'd say 1 or 2 but I'm not sure

OpenStudy (midhun.madhu1987):

If Discriminant = 0 , then the equations will have 2 real and equal roots. If Discriminant > 0 , then the equations will have 2 real roots If Discriminant < 0 , then the equations will have 2 imaginary roots

OpenStudy (anonymous):

So it's 2?? o.o

OpenStudy (midhun.madhu1987):

yes.. and both will be EQUAL

OpenStudy (anonymous):

lol that is like having one

OpenStudy (anonymous):

if \(b^2-4ac=0\) then it is a perfect square there is one zero it is \(-\frac{b}{2a}\)

OpenStudy (midhun.madhu1987):

But the Theory of Quadratic Equations states that it has 2 Equal roots..

OpenStudy (anonymous):

So it's not 2 i'm lost

OpenStudy (anonymous):

it is one not two one

OpenStudy (anonymous):

that one is often called a "repeated root" but it is one number

OpenStudy (anonymous):

Alright, thank you.

OpenStudy (anonymous):

for example if \((x-3)^2=0\) then \(x=3\) is the only solution

OpenStudy (anonymous):

yw

OpenStudy (midhun.madhu1987):

The answer is 1 only.. but this is just an added information.. since the theory states there will be 2 roots... :)